You are performing a cohort study. If the probability of developing disease in the exposed group is .05 for the study duration, then if you sample (randomly) 500 exposed people, how many do you expect to develop the disease?
Answer by stanbon(75887)

You are performing a cohort study. If the probability of developing disease in the exposed group is .05 for the study duration, then if you sample (randomly) 500 exposed people, how many do you expect to develop the disease?
----
Ans: 0.05*500 = (5/100)500 = 25
-----------
Cheers,
Stan H.
--------